# Natural Language Processing for Dialects of a Language: A Survey

### v2

@@ 5.4.1. Use of existing parsers
(Eggleston and O’Connor, 2022) parse tweets in Standard American English and African-American English and use it to analyse social attributes of an entity, as per sentiment expressed in the tweets. (Kåsen et al., 2022) create a tree bank of sentences in the Bokmål variety of Norwegian. They present their results on the UUParser, an existing parser for Norwegian. (Roy et al., 2020) present an analysis using Stanford parser and Allen NLP parser on parsing of news headlines in Indian English. (Scannell, 2020) create a treebank for Manx Gaellic and compare the performance of existing classifiers with Irish Gaellic and Scottish Gaellic.
@@ 5.4.2. Adaptation of an existing parser
:  (Chiang et al., 2006) show how parsing of Arabic dialects can be done by a sentence transduction approach. This approach parses the standardised version of a dialectic sentence, and then link it to the original sentence. The standardisation is achieved using transduction, akin to n-gram decoding. However, (Blodgett et al., 2018) uses neural networks and presents an approach to dependency parsing for African-American English. This approach uses two neural parsers, which are modified with the word embeddings used for initialisation. The word embeddings are trained on the standard and the dialect-specific datasets. Further, (Wang et al., 2017) create a dependency parser for Singaporean English. This approach uses a base parser for standard English and stacks it with a series of BiLSTM layers known as the ‘feature stack’ to extract relevant features, and an MLP with an output layer help produce dependency-parsed output. (Zhao et al., 2020) use a treebank of learner English sentences labeled with POS tags and dependency information. They propose a factorisation-based parser that first predicts nodes followed by edges in a dependency parse. (Dou et al., 2023) evaluates various parsers designed for converting text to SQL, focusing on a multilingual benchmark that covers dialects from seven different languages. This research is significant for its emphasis on semantic parsing, differing from the aforementioned dependency parsing works.
@@ 5.4.3. Development of a new parser
(Bowers et al., 2017) present a finite-state machine-based parser for the endangered Odawa dialect of Ojibwe spoken in Canada and neighbouring countries. This approach uses a phonological module composed with a morphological module where morphological strings are modified by the phonology until they match surface forms of the language.
@@ 5.5. NLU Benchmarks
Finally, NLU benchmarks such as GLUE have become popular in NLP research. They provide a set of datasets for several NLP tasks. This section discusses dialectic studies evaluated on benchmarks for NLU tasks. (Ziems et al., 2022) show a drop in performance on 7 GLUE tasks including SST-2, SST-B. Drop in performance is observed. For SST-2 it is 1.5-2% drop using fine-tuned RoBERTa.
@@
(Dacon et al., 2022) work with the African-American English. They first propose CodeSwitch, a rule-based method of perturbing a sentence from standard American English to African-American English. They create perturbed versions of the dataset using CodeSwitch and manually evaluate it. They finally evaluate their method on NLI. In order to do so, they use adversarial learning that ensures that the predicted label is same if either the SAE or AFE sentences are provided as the input. They refer to this as disentanglement of language style.  (Tan et al., 2020) present base-inflection encoding: a mechanism to inject dialectic information into the encoder. They show that their encoding algorithm improves the performance for Vernacular African-American English for SQUAD and MNLI tasks.
@@
(Held et al., 2023) model natural language understanding for dialects as a dialect adaptation task. Using Multi-VALUE, they create African-American English variations of the GLUE benchmark (which is primarily written in Standard American English). Following that, they adapt a model pre-trained on Standard American English. To do so, they use: (a) a contrastive loss to ensure the representation of a standard sentence and its dialectic version is as close as possible; (b) a morphosyntactic loss based on word-level alignment between the standard and dialectic sentences. Their results show improved robustness on 4 dialects based on the GLUE benchmark.
@@
A recent work by (Xiao et al., 2023) shows how low-rank adaptation (LoRA) can use linguistic knowledge of dialects to improve zero-shot performance on NLU tasks. They integrate hypernetworks with LoRA adapters for dialect adaptation. Experts encode linguistic information in the form of feature vectors. A hypernetwork then learns to generate adapter weights for LoRA from the feature vectors. They demonstrate the impact of their fine-tuning approach on several GLUE tasks such as MNLI, RTE and so on. The dataset consists of variants of the GLUE benchmark for five dialects: African American Vernacular English (AAVE), Indian English (IndE), Nigerian English (NgE), Colloquial Singaporean English (CollSgE), and Chicano English (ChcE). Similarly, (Liu et al., 2023) use dynamic aggregation of linguistic rules to adapt LLMs to multiple dialects. They first create a synthetic dataset of linguistic transformations using LLM probing. Following that, they train a set of feature adapters to generalise across multiple dialects of interest. They present their evaluation of multiple dialects of English.
@@ 5.6. Others
(Erdmann et al., 2018) investigate how word embeddings trained on dialect-specific or mixed-dialect corpora perform. In their experiments for text in dialects of Arabic, they show how dialect-specific embeddings can be helpful for dictionary induction. Dictionary induction here refers to alignment tables between dialects of a language. (Demszky et al., 2021) report models that predict dialect features using minimal pairs that represent linguistic properties of dialects. They do so for Indian English.
@@
width = colspec = Q[285]Q[224]Q[490], row1 = c, row2 = c, row5 = c, cell21 = c=30.941, cell51 = c=30.941, hline1-3,5-7 = -, Paper & Dialect Approach
@@
Table 7. Representative Examples for sequence-to-sequence NLP tasks (except machine translation).
@@ 6. Natural Language Generation (NLG)
The previous section showed that NLU for dialects has been primarily focused on tasks like identification of dialects, and sentiment analysis. We now present approaches in NLG. NLG deals with sequence-to-sequence (seq2seq) tasks in NLP which take a sequence as input and produce a sequence. Some examples of such problems are summarisation, question answering and machine translation, and are described in Table 7. While the situation in the case of NLU was already dire, our survey indicates that for NLG, it is even worse. We will now discuss NLP approaches that deal with dialects of a language in the context of seq2seq problems.
@@
    Making evaluation metrics dialect-aware(Sun et al., 2023) state that metrics used to measure text generation may penalise outputs in certain dialects. They propose a metric named NANO which allows perturbations in the generated output. They show that models pretrained with NANO as the metric can be helpful for dialect-robustness.
@@
    Creating dialectic variants of datasets for benchmarking(Ziems et al., 2023) present Multi-VALUE, a library that creates dialectic variations of datasets based on a set of manually created rules. They create variants of benchmark datasets, and evaluate the variants for several seq2seq tasks including machine translation, question answering and so on. The models for evaluation are based on modern LLMs such as BERT, ROBERTA, BART and T5. The library provides a useful resource as well as insights for dialect-aware benchmarking in the future.
@@ 6.1. Summarisation
Past work in summarisation, although limited, states that dialect labels may not be explicitly necessary. However, a review of Arabic text summarisation by (Elsaid et al., 2022) state the use of “dialect period frameworks” to incorporate semantic information about dialects. In the case of multi-document summarisation, clustering of sentences in the input set is a predominant paradigm. Two such works are noteworthy:
@@
    (Olabisi et al., 2022) analyse the diversity of dialects in multi-document summarisation of social media posts. They present a dataset that contains summaries of a collection tweets written in three dialects: African-American (AA) English, Hispanic English, and White English. They use extractive summarisation using LONGFORMER-EXT and abstractive summarisation using BART and T5. In order to bring diversity-awareness in summarisation, they create automatic clusters of input documents based on semantic attributes. They follow a 2-stage approach where the summarisers are separately applied, and the resultant outputs are combined again using a summariser.
@@
    (Keswani and Celis, 2021) examine the role of dialect diversity on multi-tweet summarisation. They use a variety of summarisers: typical traditional summarisers like TF-IDF, TextRank and LexRank, and SummaRunner (a neural summariser that treats summarisation as a sequence classifiation task). They create a control set: a subset of sentences that represent different dialects in the set of sentences. They introduce a bias mitigation procedure that introduces dialect-awareness in summaries using a parameter that is weighted to increase the score of dialect-diverse sentences in the dialect set.
@@ 6.2.1. MT between dialects of the same language
The primary goal of inter-dialect translation is the dissemination of information available between a standard dialect and a non-standard one. In this context, the following works are relevant. Mapping from less used dialects to their most common versions is called dialect normalization. One such work by Kuparinen et al. (2023) provides a dialect normalization dataset in Swiss German, Slovene, Finnish, and Norwegian. (Bouamor et al., 2014) present a multi-dialectal dataset for various dialects of Arabic.
@@
Character level modeling:  (Abe et al., 2018) conduct Japanese dialect translation where they use NMT to translate from dialect to standard Japanese using character RNN trained on small datasets collected as a part of their work.  (Honnet et al., 2018) additionally suggest that normalization is an important aspect for translating between Swiss German dialects, which is achievable via character-level models. Kuparinen et al. (2023) further show that sliding-window-based approaches are useful since dialect translation does not need the entire sentence-level context.
Perturbation-based regularization:  (Liu et al., 2022) present a seq2seq approach for machine translation of Singaporean English to standard English. They use word perturbation and sentence perturbation to prevent overfitting of lexical features.  (Maurya et al., 2023) used a similar approach for Indian dialects.
Harnessing Linguistic Features: Erdmann et al. (2017) focus on translation among Arabic dialects in a low-resource setting where they supplement small parallel corpora with morpho-syntactic information injected into the model for machine translation. In general, incorporating linguistic features into the MT framework is known to significantly boost translation quality in low-resource settings (Chakrabarty et al., 2022, 2020). Especially, pre-training by leveraging linguistic features, as done by Chakrabarty et al. (2022), should beneficial for dialectic translation which is typically a low-resource problem.
Leveraging ASR:  (Plüss et al., 2023) report a dataset of speech transcripts that map Standard German sentences to Swiss German sentences. They use XLS-R to train a system for automatic speech recognition and report a high average BLEU of 74.7. Their model reports significant gains over two other Swiss-German ASR test sets, indicating the efficacy of this corpus.
Code-mixed training: (Lu et al., 2022) use XLM for Translation between Hokkein-Mandarin code-mixed text. They observe that continuous training with code-mixed data enables monolingual language models to provide better performance when applied to code-mixed tasks.
Data Creation for MT Between Dialects:  Zbib et al. (2012) and Meftouh et al. (2015) also focus on multi-dialect MT data collection for Arabic, which is, once again, to be noted as one of the most studied languages for dialects. Xu et al. (2015) use a Hidden Markov-based model to create word alignment between dialects of Chinese: Mainland Chinese, Hong Kong Chinese, and Taiwan Chinese. The outcome is a monolingual corpus that contains corresponding words used in the three dialects. Their approach was shown to be effective for three different alignment mapping cases. Rather than use word alignment (Hassani, 2017) work on Kurdish dialectic MT using dictionaries that show that having limited to no parallel corpora is not a huge barrier for inter-dialect translation.
@@ 6.2.2. MT between dialects and another language
The second category, involving the harder challenge of machine translation between a dialect and another language, has received far more attention. We cover notable works in this section.
@@
Dialect pivoting: An early work in this regard is by (Paul et al., 2011). They present a pivot-based MT approach for the translation of four dialects of Japanese, namely Kumamoto, Kyoto, Okinawa, and Osaka. In order to map sentences across dialects, they use a character-based generative graphical model. They then translate the dialects into four Indo-European languages, using standardised Japanese as the pivot language.  (Jeblee et al., 2014) focus on using modern standard Arabic as a pivot when translating from English to the Egyptian Arabic dialect.
@@
Unsupervised segmentation: Different from Abe et al. (2011) who focus on characters,  (Al-Mannai et al., 2014; Salloum and Habash, 2022a) work on Arabic dialectic translation which shows that unsupervised word segmentation helps for translation into English.
@@
Evaluating existing translators on dialectic datasets(Kantharuban et al., 2023) show the performance of MT between English and dialects of seven languages. Using state-of-the-art MT systems such as Google NMT and Meta NLLB, they evaluate MT in both directions (to and from English). They report a drastic drop in BLEU for dialects of German, Portuguese and Bengali.
@@
Using inferred dialectic labels to guide translation(Sun et al., 2023) add language-dialect information as predicted by a language identifier as an input when training an MT system. They further improve the metrics for robust evaluation of text generation systems for different languages and dialects. They report their results on several dialects of English, Chinese, Portuguese and so on. They use few-shot prompting to create semantic perturbations to train T5. The results show that dialect-awareness improves the performance of translation. (Shapiro and Duh, 2019) explore a multidialect system and identify when dialectic identification is useful. (Tahssin et al., 2020) focus on dialect identification itself using AraBERT models. (Salloum et al., 2014) SMT work that focuses on sentence-level dialect identification for MT model selection where the MT model is optimized for that dialect.
@@
Learning through exemplars: Few-shot learning involves the use of examples in a prompt to guide the generation through a language model. (Riley et al., 2023) present a few-shot machine translation approach for translation between English and two variants of Portuguese and Chinese. The parallel corpus is manually created by native speakers. The exemplars used in the dataset are from the specific dialect in order to obtain translations of English sentences.
@@
User-generated Content: User-generated content is often mistranslated on social media, especially, for low-resource languages like dialectical Arabic. (Saadany et al., 2022) train a Transformer to translate from Dialectical Arabic to English where they focus on challenges in translation of user-generated content, and propose a sentiment-aware evaluation metric for translation. They discuss results on multiple test sets, including a hand-crafted test set, and analyze the performance of a semi-supervised approach compared to a baseline NMT system, a pivoting-based system, and Google Translate.
@@
Use of multiple translation models: Translation models that translate between the standard version and a dialect can assist machine translation. (Kumar et al., 2021) show an approach for MT from English to Ukrainian, Belarusian, Nynorsk, and Arabic dialects. They use two models: a dialect-to-standard translation model, and a standard source-to-target language translation model.
Data Creation for MT of Dialect to Another Language: (Hassan et al., 2017) explore synthetic data creation using word pairs between dialects based on embeddings. They take seed data, transform it into its dialectic variant and now have a dialectic parallel corpus. Similarly, (Almansor and Al-Ani, 2017) focus on using monolingual data and tiny parallel corpora in conjunction between cross-dialectic embeddings to improve MT between dialects.  Sajjad et al. (2020) take dialect MT evaluation further and focus on multi-domain coarse-grained analysis of dialects of Arabic via their AraBench benchmark. Hamed et al. (2022) propose an Arabic-English code-switched speech translation dataset which represents a practical use case since a vast majority of dialects are often spoken. There is a significant dearth of code-mixed datasets and recommend researchers to focus on the same. Alkheder et al. (2023), recognizing the increasing usage of Arabic in several regions of Turkey, expand the MADAR corpus (Bouamor et al., 2018a) to enable benchmarking of translation between Arabic and Turkish.

### v4

@@ 5.4.1. Use of existing parsers
Eggleston and O’Connor (2022) parse tweets in Standard American English and African-American English and use it to analyse social attributes of an entity, as per sentiment expressed in the tweets. Kåsen et al. (2022) create a tree bank of sentences in the Bokmål variety of Norwegian dialects. They present their results on the UUParser, an existing parser for Norwegian. Roy et al. (2020) present an analysis using Stanford parser and Allen NLP parser on parsing of news headlines in Indian English. Scannell (2020) create a treebank for Manx Gaelic and compare the performance of existing classifiers with Irish Gaelic and Scottish Gaelic.
@@ 5.4.2. Adaptation of an existing parser
Chiang et al. (2006) show how parsing of Arabic dialects can be done by a sentence transduction approach. This approach parses the standardised version of a dialectal sentence, and then links it to the original sentence. The standardisation is achieved using transduction, akin to n-gram decoding. However, Blodgett et al. (2018) use neural networks and present an approach to dependency parsing for African-American English. This approach uses two neural parsers, which are modified with the word embeddings used for initialisation. The word embeddings are trained on the standard and the dialect-specific datasets. Further, Wang et al. (2017) create a dependency parser for Singaporean English. This approach uses a base parser for standard English and stacks it with a series of BiLSTM layers known as the ‘feature stack’ to extract relevant features, and an MLP with an output layer to help produce dependency-parsed output. Zhao et al. (2020) use a treebank of learner English sentences labeled with POS tags and dependency information. They propose a factorisation-based parser that first predicts nodes followed by edges in a dependency parse. Dou et al. (2023) evaluates various parsers designed for converting text to SQL, focusing on a multilingual benchmark that covers dialects from seven different languages. This research is significant for its emphasis on semantic parsing, differing from the aforementioned dependency parsing works.
@@ 5.4.3. Development of a new parser
: Vaillant (2008) propose a rule-based approach to construct a common syntactic description for a group of Creoles from Haiti, Guadeloupe, Martinique and French Guiana. Bowers et al. (2017) present a finite-state machine-based parser for the endangered Odawa dialect of Ojibwe spoken in Canada and northeastern United States. This approach uses a phonological module composed of a morphological module where morphological strings are modified by the phonology until they match surface forms of the language.
@@ 5.5. NLU Benchmarks
Finally, benchmarks such as GLUE, which provide datasets for NLP tasks like semantic textual similarity prediction (STS-B), sentiment classification (using the Stanford sentiment treebank (SST-2)), natural language inference (NLI), textual entailment and so on, are an important part of language model evaluation pipeline. Ziems et al. (2022) show a drop in performance on 7 GLUE tasks including SST-2, STS-B, when tested on dialectal English variations of the original Standard American English version. For example, for SST-2, there is a 1.5-2% drop using fine-tuned RoBERTa.
@@
Dacon et al. (2022) work with African-American English. They first propose CodeSwitch, a rule-based method of perturbing a sentence from Standard American English (SAE) to African-American English (AAE). They create perturbed versions of the dataset using CodeSwitch and manually evaluate it. They finally evaluate their method on NLI. In order to do so, they use adversarial learning that ensures that the predicted label is the same if either the SAE or AAE sentences are provided as the input. They refer to this as a disentanglement of language style.  Tan et al. (2020) present base-inflection encoding: a mechanism to inject dialectal information into the encoder. They show that their encoding algorithm improves the performance of Vernacular African-American English for SQUAD and MNLI tasks.
@@
Held et al. (2023) model natural language understanding for dialects as a dialect adaptation task. Using Multi-VALUE, they create African-American English variations of the GLUE benchmark (which is primarily written in Standard American English). Following that, they adapt a model pre-trained on Standard American English. To do so, they use: (a) a contrastive loss to ensure the representation of a standard sentence and its dialectal version is as close as possible; (b) a morphosyntactic loss based on word-level alignment between the standard and dialectal sentences. Their results show improved robustness on 4 dialects based on the GLUE benchmark. Vidal-Gorène et al. (2024) provide a benchmark on lemmatization, POS-tagging, and morphological analysis for four Armenian varieties- Classical, Modern Eastern, Modern Western, and the under-documented Getashen dialect. They compare traditional RNN models, multilingual encoders, and large language models using supervised, transfer learning, and zero/few-shot learning approaches, and show how RNNs are strong at POS tagging, but LLMs handle unseen dialectal variations.
@@
A recent work by Xiao et al. (2023) shows how low-rank adapters Low-Rank Adapters (LoRA) (a parameter-efficient fine-tuning or PEFT technique that allows fine-tuning LLMs faster by storing weight updates instead of updating all weights) can use linguistic knowledge of dialects to improve zero-shot performance on NLU tasks. They integrate hypernetworks with LoRA adapters for dialect adaptation. Experts encode linguistic information in the form of feature vectors. A hypernetwork then learns to generate adapter weights for LoRA from the feature vectors. They demonstrate the impact of their fine-tuning approach on several GLUE tasks such as MNLI, RTE and so on. The dataset consists of variants of the GLUE benchmark for five dialects: African American Vernacular English (AAVE), Indian English (IndE), Nigerian English (NgE), Colloquial Singaporean English (CollSgE), and Chicano English (ChcE). Similarly, Liu et al. (2023) use dynamic aggregation of linguistic rules to adapt LLMs to multiple dialects. They first create a synthetic dataset of linguistic transformations using LLM probing. Following that, they train a set of feature adapters to generalise across multiple dialects of interest. They present their evaluation of multiple dialects of English.
DIALECTBENCH (Faisal et al., 2024) is a large-scale benchmark covering 10 NLP tasks focusing on 281 language varieties. Their evaluation shows substantial disparities in performance between the standard and non-standard language varieties, while also identifying language clusters with large performance divergence across tasks. Most recently, the VarDial 2024 evaluation campaign (Chifu et al., 2024) released dataset on the choice of plausible alternatives (COPA) task focusing on three micro-dialects namely, Cerkno dialect of Slovenian, Chakavian dialect of Croatian, and the Torlak dialect which is spoken across Serbia, Macedonia, and Bulgaria. This task requires a computational model to select one of two candidate statements which is more likely to be the cause or effect of a given premise statement. Collectively, training and test datasets from VarDial evaluation campaigns ($2014$ $2024$) organised over the years should act as a good benchmark for LLM evaluation of dialects.
@@ 5.6. Others
Erdmann et al. (2018) investigate how word embeddings trained on dialect-specific or mixed-dialect corpora perform. In their experiments for text in dialects of Arabic, they show how dialect-specific embeddings can be helpful for dictionary induction. Dictionary induction here refers to alignment tables between dialects of a language. Demszky et al. (2021) report models that predict dialect features using minimal pairs that represent linguistic properties of dialects. They do so for Indian English.
@@
width = colspec = Q[285]Q[224]Q[490], row1 = c, row2 = c, row5 = c, row8 = c, row11 = c, cell21 = c=30.941, cell51 = c=30.941, cell81 = c=30.941, cell111 = c=30.941, hline1-3,5-6, 8-9, 11-13 = -, Paper & Dialect Approach
@@
Machine translation
 (Alam et al., 2024) 891 variations Contrastive dialect benchmark for MT evaluation
 (Riley et al., 2023) Portuguese and Chinese Dialects & English Few-shot approach for MT
Dialogue systems
 (Zhan et al., 2023) Dialectal norms in Chinese culture Benchmark dataset
 (Artemova et al., 2024)
German dialects Perturbation-based evaluation for dialects
@@
Table 7. Representative Examples for sequence-to-sequence NLP tasks.
@@ 6. Natural Language Generation (NLG)
The previous section showed that NLU for dialects has primarily focused on tasks like identification of dialects and sentiment analysis. We now present approaches in NLG. NLG deals with sequence-to-sequence (seq2seq) tasks in NLP, which take a sequence as input and produce a sequence. Challenges in the presence of dialects in a generation task can differ significantly given the task. The data and evaluation methods can be different for tasks, especially where dialectal text is being generated. Some examples of such problems are summarisation, question answering and machine translation, and are described in Table 7. While the situation in the case of NLU was already dire, our survey indicates that for NLG, it is even worse. We will now discuss NLP approaches that deal with dialects of a language in the context of seq2seq problems.
@@
    Making evaluation metrics dialect-awareSun et al. (2023) state that metrics used to measure text generation may penalise outputs in certain dialects. They propose a metric named NANO, which allows perturbations in the generated output. They show that models pretrained with NANO as the metric can be helpful for dialect-robustness.
@@
    Creating dialectal variants of datasets for benchmarkingZiems et al. (2023) present Multi-VALUE, a library that creates dialectal variations of datasets based on a set of manually created rules. They create variants of benchmark datasets, and evaluate the variants for several seq2seq tasks including machine translation, question answering and so on. The models for evaluation are based on modern LLMs such as BERT, ROBERTA, BART and T5. The library provides a useful resource as well as insights for dialect-aware benchmarking in the future.
@@ 6.1. Summarisation
Past work in summarisation, although limited, states that dialect labels may not be explicitly necessary. However, a review of Arabic text summarisation by Elsaid et al. (2022) state the use of “dialect period frameworks” to incorporate semantic information about dialects. In the case of multi-document summarisation, clustering of sentences in the input set is a predominant paradigm. Two such works are noteworthy:
@@
    Olabisi et al. (2022) analyse the diversity of dialects in multi-document summarisation of social media posts. They present a dataset that contains summaries of a collection of tweets written in three dialects: African-American English, Hispanic English, and White English. They use extractive summarisation using LONGFORMER-EXT and abstractive summarisation using BART and T5. In order to bring diversity-awareness in summarisation, they create automatic clusters of input documents based on semantic attributes. They follow a 2-stage approach where the summarisers are separately applied, and the resultant outputs are combined again using a summariser.
@@
    Keswani and Celis (2021) examine the role of dialect diversity on multi-tweet summarisation. They use a variety of summarisers: typical traditional summarisers like TF-IDF, TextRank and LexRank, and SummaRunner (a neural summariser that treats summarisation as a sequence classifiation task). They create a control set: a subset of sentences that represent different dialects in the set of sentences. They introduce a bias mitigation procedure that introduces dialect-awareness in summaries using a parameter that is weighted to increase the score of dialect-diverse sentences in the dialect set.
@@ 6.2.1. MT between dialects of the same language
The primary goal of inter-dialect translation is the dissemination of information available between a standard dialect and a non-standard one. In this context, the following works are relevant. Mapping from less used dialects to their most common versions is called dialect normalisation. One such work by Kuparinen et al. (2023) provides a dialect normalisation dataset in Swiss German, Slovene, Finnish, and Norwegian. Bouamor et al. (2014) present a multi-dialectal dataset for various dialects of Arabic.
@@
Character level modeling:  Abe et al. (2018) conduct Japanese dialect translation where they use NMT to translate from dialect to standard Japanese using character RNN trained on small datasets collected as a part of their work.  Honnet et al. (2018) additionally suggest that normalisation is an important aspect for translating between Swiss German dialects, which is achievable via character-level models. Kuparinen et al. (2023) further show that sliding-window-based approaches are useful since dialect translation does not need the entire sentence-level context.
Perturbation-based regularisationLiu et al. (2022) present a seq2seq approach for machine translation of Singaporean English to standard English. They use word perturbation and sentence perturbation to prevent overfitting of lexical features. Maurya et al. (2023) used a similar approach for Indian dialects.
Harnessing Linguistic Features: Erdmann et al. (2017) focus on translation among Arabic dialects in a low-resource setting where they supplement small parallel corpora with morpho-syntactic information injected into the model for machine translation. In general, incorporating linguistic features into the MT framework is known to significantly boost translation quality in low-resource settings (Chakrabarty et al., [n. d.], 2020). Especially, pre-training by leveraging linguistic features, as done by Chakrabarty et al. ([n. d.]), should be beneficial for dialectal translation, which is typically a low-resource problem.
Code-mixed training: Lu et al. (2022) use XLM for Translation between Hokkien-Mandarin code-mixed text. They observe that continuous training with code-mixed data enables monolingual language models to provide better performance when applied to code-mixed tasks.
Data Creation for MT between dialects:  Zbib et al. (2012) and Meftouh et al. (2015) also focus on multi-dialect MT data collection for Arabic, which is, once again, to be noted as one of the most studied languages for dialects. Xu et al. (2015) use a Hidden Markov-based model to create word alignment between dialects of Chinese: Mainland Chinese, Hong Kong Chinese, and Taiwan Chinese. The outcome is a monolingual corpus that contains corresponding words used in the three dialects. Their approach was shown to be effective for three different alignment mapping cases. Rather than use word alignment, Hassani (2017) works on Kurdish dialectal MT using dictionaries and show that having limited to no parallel corpora is not a significant barrier for inter-dialect translation. YORULECT (Ahia et al., 2024) releases parallel text and speech corpus for four regional variants of Yorùbá, by collecting data from within native communities for each variant, and providing a benchmark for machine translation, speech recognition, and text-to-speech synthesis. Most recently, Dabre et al. (2024) worked on Kadodi, a dialect of Marathi, for which the local community was involved for parallel corpora creation, further highlighting the need for community involvement.
@@ 6.2.2. MT between dialects and another language
The second category, involving the harder challenge of machine translation between a dialect and another language, has received far more attention. We cover notable works below.
@@
Dialect pivoting: An early work in this regard is by Paul et al. (2011). They present a pivot-based MT approach for the translation of four dialects of Japanese, namely Kumamoto, Kyoto, Okinawa, and Osaka. In order to map sentences across dialects, they use a character-based generative graphical model. They then translate the dialects into four Indo-European languages, using standardised Japanese as the pivot language.  Jeblee et al. (2014) focus on using modern standard Arabic as a pivot when translating from English to the Egyptian Arabic dialect.
@@
Unsupervised segmentation: Different from Abe et al. (2011) who focus on characters,  Al-Mannai et al. (2014); Salloum and Habash (2022) work on Arabic dialectal translation, which shows that unsupervised word segmentation is just as effective if not better for translation into English.
@@
Evaluating existing translators on dialectal datasetsKantharuban et al. (2023) show the performance of MT between English and dialects of seven languages. Using state-of-the-art MT systems such as Google NMT and Meta NLLB, they evaluate MT in both directions (to and from English). They report a drastic drop in BLEU for dialects of German, Portuguese and Bengali.  De Camillis et al. (2023) train NMT systems to evaluate the performance of legal domain translation for Italian $\Leftrightarrow$ South Tyrolean German, where their models show better performance compared to Google Translate and DeepL for this niche use case. Similarly, CODET (Alam et al., 2024) is a contrastive dialectal benchmark dataset for evaluating machine translation systems focusing on $891$ variations from $12$ languages.
@@
Using inferred dialectal labels to guide translationSun et al. (2023) add language-dialect information as predicted by a language identifier as an input when training an MT system. They further improve the metrics for robust evaluation of text generation systems for different languages and dialects. They report their results on several dialects of English, Chinese, Portuguese and so on. They use few-shot prompting to create semantic perturbations to train T5. The results show that dialect-awareness improves the performance of translation. Shapiro and Duh (2019) explore a multidialect system and identify when dialectal identification is useful. Tahssin et al. (2020) focus on dialect identification itself using AraBERT models. Salloum et al. (2014) focus on sentence-level dialect identification for an SMT model selection where the model is optimised for that dialect.
@@
Learning through exemplars: Few-shot learning involves the use of examples in a prompt to guide the generation through a language model. Riley et al. (2023) present a few-shot machine translation approach for translation between English and two variants of Portuguese and Chinese. The parallel corpus is manually created by native speakers. The exemplars from the specific dialect are used to obtain translations of English sentences.
@@
User-generated content: User-generated content (UGC) is often mistranslated on social media, especially, for low-resource languages like dialectal Arabic. Saadany et al. (2022) train a Transformer to translate from dialectal Arabic to English where they focus on challenges in translation of UGC, and propose a sentiment-aware evaluation metric for translation. They discuss results on multiple test sets, including a hand-crafted test set, and analyse the performance for a semi-supervised approach compared to a baseline NMT system, a pivoting-based system, and Google Translate.
@@
Use of multiple translation models: Translation models that translate between the standard version and a dialect can assist machine translation. Kumar et al. (2021) show an approach for MT from English to Ukrainian, Belarusian, Nynorsk, and Arabic dialects. They use two models: a dialect-to-standard translation model, and a standard source-to-target language translation model.
[…]